OpenStudy (anonymous):

The midpoint of a line segment with end points as (-10, y1) and (-6, 7) is (-8, 6). What is the value of y1? -15 2 5 -1

OpenStudy (anonymous):

Can you tell me what is the sum of y coordinates there??

OpenStudy (anonymous):

uhmm idk

OpenStudy (anonymous):

Sum of y coordinates of end points I am asking..

OpenStudy (anonymous):

how do you find the endpoints?

OpenStudy (anonymous):

See, the sum is : \[y_1 + 7\] Divide it by 2, then it will become equal to 6 (y coordinate of midpoint) \[\frac{y_1 + 7}{2} = 6\] Multiply 2 both the sides and can you find \(y_1\) from here??

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

Take your time..

OpenStudy (anonymous):

\[y ^{1}+7\div 2\]

OpenStudy (anonymous):

hold on this is kinda hard

OpenStudy (anonymous):

Hard??? multiply by 2 both the sides and show me what you get??

OpenStudy (anonymous):

\[2 \times \frac{(y_1 + 7)}{2} = 2 \times 6\] Can you solve this??

OpenStudy (anonymous):

2y^1 plus 14 over 2 = 12

OpenStudy (anonymous):

Why are you not cancelling 2 with 2???

OpenStudy (anonymous):

ok i will cancel it right now

OpenStudy (anonymous):

2y^1 plus 14 = 12

OpenStudy (anonymous):

2y^1=-2

OpenStudy (anonymous):

How???

OpenStudy (anonymous):

divide 2 to both sides and i think the answer is y=1

OpenStudy (anonymous):

What are you doing??

OpenStudy (anonymous):

i'm not sure this is confusing

OpenStudy (anonymous):

Solve for LHS first.. Tell me what do you get: \[2 \times \frac{(y_1+7)}{2}\] What you get??

OpenStudy (anonymous):

You have to simply cancel 2 with 2 and answer is the remaining part.. Solve this and tell me what do you get..

OpenStudy (anonymous):

y^1 plus 7

OpenStudy (anonymous):

Yes now you are right.. \[y_1 + 7 = 6 \times 2\] Solve the RHS now: What is 6*2 = ???

OpenStudy (anonymous):

12

OpenStudy (anonymous):

So write the entire now what it becomes after solving.

OpenStudy (anonymous):

y^1 plus 7=12

OpenStudy (anonymous):

Now subtract both the sides by 7 and you will have your answer...

OpenStudy (anonymous):

y^1=5

OpenStudy (anonymous):

Yes this is your answer..

OpenStudy (anonymous):

ty very much

OpenStudy (anonymous):

Welcome dear..